The beta for stock (A) is 1.40, the beta for stock (B) is 0.90, and the beta for stock (C) is 1.00. If you are equally invested in all three, what will be your portfolio’s beta?

Group of answer choices a. 1.10 b. 1.13 c. 1.18 d. 1.20
Answer #1

Option (a) is correct

Portfolio beta is the weighted average beta of the stocks in the portfolio. Since it is an equally invested portfolio, so weights of the stocks are 1/3, 1/3, 1/3:

In the next step we will multiply the weights with the betas of the stocks to find the weighted betas of the stocks:

Stock A: 1/3 * 1.4 = 0.4666

Stock B: 1/3 * 0.9 = 0.3

Stock C: 1/3 * 1 = 0.3333

Now, we will calculate the weighted average beta by adding the weighted betas of the stocks calculated above:

Weighted average beta = 0.4666 + 0.3 + 0.3333 =  1.10